OpenStudy (anonymous):

Please help me solve this question: the number of diagonals of an n-sided polygon is given by the formula D=n/2 (n-3). A polygon has 90 diagonals. How many sides does it have?

OpenStudy (anonymous):

Solve the quadratic.

OpenStudy (anonymous):

but i seem to get stuck when i get up too 180=n^2-3n, how do i solve it, the quadratic formula won't work.

OpenStudy (anonymous):

The quadratic formula works. Also for quadratic of form x^2 + bx + c , u can try to find 2 numbers p and q that add up to b (-3 in this case) and multiply together to give c (-180 in this case). Then the factors are (x+p)(x+q).

OpenStudy (anonymous):

Here is quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] Explain me why doesn't work?

OpenStudy (anonymous):

ok then, my mistake, because when i substituted in the numbers, it turned out to be a negative number under the square root

OpenStudy (anonymous):

-3^2 - 4.-180 is positive.

OpenStudy (anonymous):

woopsy, okay soz... thanks for helping!

OpenStudy (anonymous):

ur welcome.